OpenStudy (anonymous):

(i) Show that for all values of m, the line y = mx − 3m2 touches the parabola x2 = 12y. (ii) Find the values of m for which this line passes through the point (5, 2). (iii) Hence determine the equations of the two tangents to the parabola x2 = 12y from the point (5, 2). Anyone know how to do this ?

OpenStudy (anonymous):

any point on the parabola (x^2=4*3*y) is of the form (6t,3t^2) differntiating x^2=12y we have 2x=12(dy/dx) or dy/dx =x/6 at (6t,3t^2) dy/dx =6t/6=t (slope of all the tangents passing through (6t,3t^2) on the parabola) hence eq of tangent through (6t,3t^2) is of the form then y-3t^2=t(x-6t) or the eq of tangent is y=tx +3t^2-6t^2 or y=tx -3t^2 that is to say for all real values of 't' y=tx-3t^2 is always tangent to the x^2=12y hence if t=m then y=mx-3m^2 is always tangent to the given parabola (procved)

OpenStudy (anonymous):

if y=mx-3m^2 passes through (5,2) then 2=5m-3m^2 or 3m^2-5m+2=0 or 3m^2-3m-2m+2=0 or 3m(m-1)-2(m-1)=0 or (m-1)(3m-2)=0 thus the vlues of m are 1 and 2/3 and for m=1 tangent is y=x-3 and for m=2/3 tangent is y=(2/3) x -4/3

OpenStudy (anonymous):

hope u understand the method rather than this solution..

OpenStudy (anonymous):

is it that u r mastering tangent and normals using derivatives..

OpenStudy (anonymous):

the first part i do not understand at all, it got me even more confused but the second part i understand

OpenStudy (anonymous):

see what is ur confusion in the first part

OpenStudy (anonymous):

i just do not understand how you got 6t, 3t^2 ?

OpenStudy (anonymous):

for the parabola x^2=4by the parametric eq can be written as x=2bt and y=bt^2 t being the parameter or in other words any point on the above parabola is of the form (2bt,bt^2)

OpenStudy (anonymous):

thus here the parabola is x^2=12y or x^2=4(3)y that is b here is 3 hence any point on the parabola is of the form (2*3*t,3*t^2)

OpenStudy (anonymous):

hope u get this n further say what s still creating confusion

OpenStudy (anonymous):

are u there...